There are three industrial firms in Happy Valley.

Firm Initial Pollution Level Cost of Reducing Pollution by 1 Unit

A 30 $20

B 40 $30

C 20 $10

The government wants to reduce pollution to 60 units, so it gives each firm 20 tradable

pollution permits.

a) Who sells permits and how many do they sell? Who buys permits and how many do

they buy? Briefly explain why the sellers and buyers are each willing to do so. What is

the total cost of pollution reduction in this situation?

b) How much higher would the costs of pollution reduction be if the permits could not be

traded?


1
Expert's answer
2021-11-24T12:31:23-0500

Solution:

a.). The cost of reducing pollution is lowest for Firm C, which means that firm C would reduce pollution and sale 20 permits to firm B. The cost of reducing pollution is maximum for firm B, so firm B would buy 20 permits to firm C.

The sellers and buyers are each willing to do so in order to maximize on pollution costs.


The total cost of pollution reduction is $600. This is calculated as follows:

Firm B: Buys 20 permits

Remaining pollution reduction = 40 – 40 = 0

Total pollution reduction cost = 30 "\times" 0 = 0


Firm C: Sells 20 permits

Remaining pollution reduction = 20 – 10 = 10

Total pollution reduction cost = 10 "\times" 20 = 200


Firm A: Neither buys nor sells

Remaining pollution reduction = 30 – 10 = 20

Total pollution reduction cost = 20 "\times" 20 = 400


Total pollution reduction cost = 0 + 200 + 400 = $600


b.). If permits could not be traded. The total cost of pollution reduction is $800. This is calculated as follows:

Firm B:

Remaining pollution reduction = 40 – 20 = 20

Total pollution reduction cost = 30 "\times" 20 = 600


Firm C:

Remaining pollution reduction = 20 – 20 = 0

Total pollution reduction cost = 10 "\times" 0 = 0


Firm A:

Remaining pollution reduction = 30 – 20 = 10

Total pollution reduction cost = 20 "\times" 10 = 200


Total pollution reduction cost = 600 + 0 + 200 = $800